5) twhat ia the epusalent CA) 12yn CB)1.09
Answer #1

The circuit consists of three resistors connected in parallel

Theirfore equalent resistance of the circuit is given by the formula

1/R = 1/R1+1/R2+1/R3

1/R = 1/2+1/6+1/4

1/R = 0.5+0.167+0.25

1/R = 0.916

R = 1/0.916

R = 1.09

Option (b) is correct